OpenStudy (anonymous):

Factor the expression 64a^ 2- 16ab + b^ 2 into a product of binomials

OpenStudy (anonymous):

you are supposed to recognize this as \[(a-b)^2=a^2-2ab+b^2\] with \[a=4a,b=b\]

jimthompson5910 (jim_thompson5910):

64a^ 2- 16ab + b^ 2 64a^ 2 - 8ab - 8ab + b^ 2 (64a^ 2 - 8ab) +(-8ab + b^2) 8a(8a - b) +(-8ab + b^2) 8a(8a - b) - b(8a - b) (8a - b)(8a - b) (8a - b)^2 So \[\large 64a^ 2- 16ab + b^ 2 = (8a - b)^2\]